OpenStudy (anonymous):

4(2x+1)-12x<4x-84 I worked it out and got x>1. by the way the > before the x is supposed to be underlined also the < before the 4x. I cant find that symbol though so couldnt write it that way.

OpenStudy (anonymous):

4(2x+1)-12x<4x-84 8x+4-12x<4x-84 -4x+4<4x-84 4<8x-84 88<8x 11<x

OpenStudy (anonymous):

Did you do it with the underlined <'s? I dont know what that symbols called i forget.

OpenStudy (anonymous):

The less than or greater to?

OpenStudy (anonymous):

\[ \le\] this symbol.

OpenStudy (anonymous):

Yup, I solved it using that.

OpenStudy (anonymous):

4(2x+1)-12x≤4x-84 8x+4-12x≤4x-84 -4x+4≤4x-84 4≤8x-84 88≤8x 11≤x

OpenStudy (anonymous):

dont u mean x\[ \ge\]11

OpenStudy (anonymous):

*x\[ \ge\] 11

OpenStudy (anonymous):

and the greph would be opened circle right?

OpenStudy (anonymous):

Yup, it'd be the same as x≥11. The graph would look like this.